OpenStudy (anonymous):

help please, I don't fully understand the concept of infinity:

OpenStudy (anonymous):

given: \[w=\infty\] what is:\[Z= \sqrt{R ^{2}+(wL-\frac{ 1 }{ wC})^{2}}\]

OpenStudy (anonymous):

First, the short answer is that Z is infinite as well. Where did you come across this question?

OpenStudy (anonymous):

Because in mathematics, treating "infinity" like it is some number is actually considered really sloppy. In actuality, INFINITY is not a number--it's a concept. Basically, the best way to restate your question is to say that the limit of Z as w goes to infinity is equal to positive infinity.... Is this helpful??

OpenStudy (anonymous):

@Splash_Dance sorry for the late reply, I got this is a physics formula for total impedance Z, we had to evaluate it for w= 0 and w= infinity. Don't really understand how Z will be equal to infinity..

OpenStudy (anonymous):

First, I should clarify something: I absolutely HATE physics. Second, Physicists are weird: They do things that make Mathematicians cringe (like plugging in "infinity" into a formula). Nevertheless, if what you're doing is finding the limit of "Z" as "w" goes to infinity, then "Z" definitely goes to infinity as well. But, to be frank, I can't help you with the physical intuition because--as I indicated earlier--I REALLY hate physics. What part of the problem are you confused about??

OpenStudy (anonymous):

we have the these two terms wL and 1/wC: what happens when we multiplying them with infinity gives infinity?

OpenStudy (anonymous):

Don't worry about all that--here, give me a second and I'll rewrite the expression...

OpenStudy (anonymous):

ok thank you very much! :)

OpenStudy (anonymous):

Ok, so what you're asking is: \[\lim_{w \rightarrow \infty} Z = \lim_{w \rightarrow \infty} \sqrt{R ^{2} + (wL - \frac{ 1 }{ wC })^{2}}\] \[= \sqrt{\lim_{w \rightarrow \infty}( R ^{2} + (wL - \frac{ 1 }{ wC })^{2}})\] \[=\sqrt{R ^{2} + \lim_{w \rightarrow \infty} (wL - \frac{ 1 }{ wC })^{2}}\] Now, when w goes to infinity, the term "1/(wC)" goes to zero--i.e. you can ignore it. And all that other stuff--the square root, R-squared, the square-ing, etc.--none of that matters. [Well, actually, the sign of "L" would've mattered but the squaring ensures that it's positive.] Quick question: Is this helping?

OpenStudy (anonymous):

ok I understand the derivation/simplifications and so as w approaches infinity, (wL)^2 is equal to infinity? and it is helping a lot

OpenStudy (anonymous):

Yeah, pretty much. Except you'd probably want to say it more like, "As w goes to infinity, Z increases without bound". Any other questions?

OpenStudy (anonymous):

nope, thank you very much! ^^